- y = A cos (4x) + B sin (4x)


1
Expert's answer
2021-02-24T11:48:20-0500

Given differential equation is-

"y=Acos4x+Bsin4x"


Differentiate above equation w.r.t. 'x'

"\Rightarrow \dfrac{dy}{dx}=-4Asin4x+4Bcos4x"

Again differentiating w.r.t. 'x'

"\dfrac{d^2y}{dx^2}=-16Acos4x-16Bsin4x"


"\Rightarrow \dfrac{d^2y}{dx^2}=-16(Acos4x+Bsin4x)"


"\Rightarrow \dfrac{d^2y}{dx^2}=-16(y)"


"\Rightarrow \dfrac{d^2y}{dx^2}+16y=0"